# Implicit High-Order Moment Tensor Estimation and Learning Latent Variable Models

### v1

@@ 4.2 Learning Positive Linear Combinations of ReLU Activations
Lemma 4.8 (namely, the definition of the $c_{n}$’s) implies that $T_{n}$ is zero, if $n>1$ and $n$ odd. Thus, it suffices to consider the Hermite coefficients of order $n=1$ or $n$ even. We claim that, for $n=1$ or $n$ even, we can apply Proposition 1.7 to efficiently approximate $T_{n}$. Specifically, we will show how to efficiently approximate the inner product $\langle T_{n},H_{n}(x)\rangle$ for $x\sim N(0,I)$. Given that such an efficient computation is possible, the hypothesis of our learning algorithm will be an approximation to the low-degree Hermite expansion of $F$ of the function
$\widetilde{H}(x)=\sum_{n=0}^{n_{0}}\widetilde{\langle T_{n},H_{n}(x)\rangle}\;,$ (6)
where $n_{0}=\mathrm{poly}(1/\epsilon)$. Note that if $H(x)=\sum_{n=0}^{n_{0}}\langle T_{n},H_{n}(x)\rangle$ is the (exact) low-degree Hermite expansion of $F$ up to degree $n_{0}$, them $\mathbf{E}_{x\sim N(0,I)}[(H(x)-F(x))^{2}]\leq(\epsilon/2)^{2}$. So for our purposes it suffices to approximate each term in the sum up to error $\epsilon/(2n_{0})=\mathrm{poly}(\epsilon)$.
We will show how to obtain an evaluation oracle for $\widetilde{H}$ with sample and computational complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$. To do this, we will apply Proposition 1.7 to obtain an approximation $\widetilde{\langle T_{n},H_{n}(x)\rangle}$ of $\langle T_{n},H_{n}(x)\rangle$ for all $n\leq n_{0}$ (excluding odd values greater than one). It turns out that the sample size and running time are dominated by the case $n=n_{0}$.
We proceed to apply Proposition 1.7. The parameter $m$ in the statement of the proposition will be set to $m:=n$, for $n\leq n_{0}$. The tensor $M_{t}$, for $t\leq 2m$ with $t$ even or equal to $m$, will be set to $M_{t}:=\sum_{i=1}^{t}w_{i}v_{i}^{\otimes t}$, where $w_{i}$ and $v_{i}$ are the hidden weights and parameter vectors of the target function $F(x)$.
We exhibit a sequential tensor computation $\mathcal{S}_{t}$ with the desired properties. Specifically, $\mathcal{S}_{t}$ is a sequential tensor computation with three inputs: two vector inputs $v,u\in\mathbb{R}^{d}$ and a scalar input $y\in\mathbb{R}$. The output of $\mathcal{S}_{t}(u,v,y)$ is the order-$t$ and dimension-$d$ tensor $(1/c_{t})yH_{t}(u,v)$, where $c_{t}=\Theta(t^{-5/4})$ is the explicit constant in Lemma 4.8. Since $H_{t}(u,v)$ can be computed by a sequential tensor computation of size $O(t)$, the same holds for $(1/c_{t})yH_{t}(u,v)$.
The input distribution corresponding to these inputs is the joint distribution $\mathcal{D}:=(X,Y,F(X))$, where $X$ and $Y$ are independent standard Gaussians, $X,Y\sim N(0,I)$, and $F(X)$ is the value of the target function. Since we have sample access to labeled examples from the target function, i.e., from the distribution $(X,F(X))$, the distribution $\mathcal{D}$ is efficiently samplable. We now have that
$\mathbf{E}_{(u,v,y)\sim\mathcal{D}}\mathcal{S}_{t}(u,v,y)=\mathbf{E}_{X\sim N(0,I),Y\sim N(0,I)}[(1/c_{t})F(X)H_{t}(X,Y)]=(1/c_{t})T_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}\;,$
where the first equality holds by definition, the second equality follows from (5), and the third equality follows from Lemma 4.8 (Equation (4)).
This gives us a sequential tensor computation $\mathcal{S}_{t}$ of size $O(t)$ with expectation $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$, when $t$ is equal to $1$ or even. It remains to bound above the covariance of $\mathcal{S}_{t}$. We show the following:
Claim 4.9.
We have that $\mathbf{Cov}_{(u,v,y)\sim\mathcal{D}}[\mathcal{S}_{t}(u,v,y)]$ is bounded above by $V:=O(1)^{t}$.
Proof.
Recall that for $(u,v,y)\sim\mathcal{D}$, we have that $\mathcal{S}_{t}(u,v,y)=(1/c_{t})F(X)H_{t}(X,Y)$. By definition, we want to bound from above the variance of the random variable $\langle W,(1/c_{t})F(X)H_{t}(X,Y)\rangle=(1/c_{t})F(X)\langle W,H_{t}(X,Y)\rangle$ for any tensor $W\in(\mathbb{R}^{d})^{\otimes t}$ with $\|W\|_{2}=1$. We can write
$\mathbf{Var}[\langle W,(1/c_{t})F(X)H_{t}(X,Y)\rangle]$ $=$ $(1/c_{t})^{2}\mathbf{Var}[\langle W,F(X)H_{t}(X,Y)\rangle]$
$=$ $\Theta(t^{5/2})\mathbf{Var}[F(x)\langle W,H_{t}(X,Y)\rangle]$
$\leq$ $\Theta(t^{5/2})\,\|F\|_{4}^{2}\,\|\langle W,H_{t}(X,Y)\rangle\|_{4}^{2}$
$\leq$ $\Theta(t^{5/2})\,O(1)\,O(1)^{t}\,\|\langle W,H_{t}(X,Y)\rangle\|_{2}^{2}$
$\leq$ $\Theta(t^{5/2})\,O(1)\,O(1)^{t}\,O(1)^{t}$
$=$ $O(1)^{t}\;,$
where the second line follows from the definition of $c_{t}$ (for $t=1$ or even), and the third line uses Cauchy-Schwartz. The fourth line follows from two facts. First, for any $m\in\mathbb{Z}_{+}$, we have
$\|F(X)\|_{m}\leq\sum_{i=1}^{k}w_{i}\|\mathrm{ReLU}(v_{i}\cdot X)\|_{m}\leq\sum_{i=1}^{k}w_{i}\|v_{i}\cdot X\|_{m}=O(\sqrt{m})\sum_{i=1}^{k}w_{i}=O(\sqrt{m})\;.$
@@
Second, since $\langle W,H_{t}(X,Y)\rangle$ is a degree-$t$ polynomial over Gaussians, hypercontractivity (Fact 2.4) implies that this quantity is $O(1)^{t}\|\langle W,H_{t}(X,Y)\rangle\|_{2}^{2}$.
Finally, by Lemma 4.3 (applied for $\mu=0$), it follows that $\|\langle W,H_{t}(X,Y)\rangle\|_{2}^{2}$ is at most $O(1)^{t}$. This justifies the fifth line and completes the proof of the claim. ∎
It remains to exhibit the sequential tensor computation $\mathcal{F}_{m}$, where we recall that $m=n$. $\mathcal{F}_{n}$ is a sequential tensor computation with two input vector inputs $v,u\in\mathbb{R}^{d}$. The output of $\mathcal{F}_{n}(u,v)$ is the order-$n$ and dimension-$d$ tensor $H_{n}(u,v)$. We recall the fact that $H_{n}(u,v)$ can be computed by a sequential tensor computation of size $O(n)$.
The input distribution corresponding to these inputs is the $(X,Y)$, where $X$ and $Y$ are independent standard Gaussians, $X,Y\sim N(0,I)$. We note that $\mathcal{D}^{\prime}$ is the distribution of $Y$ (that we have sample access to) while $X$ is a single Gaussian sample (see the proposition statement). Since $H_{n}(X)=\mathbf{E}_{Y}[H_{n}(X,Y)]$ and the covariance of $H_{n}(X,Y)$ is bounded above by $O(1)^{n}$ (by Lemma 4.3), we have all the ingredients we need to apply Proposition 1.7.
By Proposition 1.7, if the total number of samples from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ is $N=\mathrm{poly}(d,k,S,V,1/\epsilon)$, where $V=O(1)^{n}$, for a Gaussian random $X$ we can approximate $\langle T_{n},H_{n}(X)\rangle$ up to expected squared error $\epsilon$ in $\mathrm{poly}(N,d)$ time.
Recall that $V=O(1)^{n}$ and $S=O(n)$. By setting $n=n_{0}=\mathrm{poly}(1/\epsilon)$ and the desired accuracy $\epsilon$ to be $\epsilon/n_{0}=\mathrm{poly}(1/\epsilon)$, we obtain estimates that suffice to evaluate the hypothesis $\widetilde{H}$.
Overall, this gives a learning algorithm with sample complexity and runtime $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$, completing the proof of Theorem 4.7. ∎
@@ 4.4 Learning Mixtures of Linear Regressions
Note that if we are given an $F$ with $B$ or $\sigma$ more than $1$, we can simply divide the $y$-values by some constant $C$, which will have the effect of dividing both $B$ and $\sigma$ by $C$. For $C$ large enough, the hypotheses of Theorem 4.19 should apply.
Proof of Theorem 4.19
@@ 5 Conclusions
In this work, we gave a general efficient algorithm to approximate higher-order moment tensors, as long as there are reasonable unbiased estimators for these moments that can be efficiently sampled. This type of tensors arise in a range of learning problems. We leveraged our general result to obtain the first polynomial-time algorithms for learning mixtures of linear regressions, learning sums of ReLUs, density estimation for mixtures of spherical Gaussians with bounded means, and parameter estimation for mixtures of spherical Gaussians under optimal separation. In all cases, our learning algorithms run in $\mathrm{poly}(d,k)$ time.

### v2

@@ 4.2 Learning Positive Linear Combinations of Non-Linear Activations
In particular, we will want to use Proposition 3.2 to efficiently approximate the inner product $\langle T_{m},H_{m}(x)\rangle$ for $x\sim N(0,I)$. Given that such an efficient computation is possible, the hypothesis of our learning algorithm will be an approximation to the low-degree Hermite expansion of $F$ of the function
$\widetilde{H}(x)=\sum_{m=0}^{n}\widetilde{\langle T_{m},H_{m}(x)\rangle}.$ (8)
In particular, so long as our approximation to $\langle T_{m},H_{m}(x)\rangle$ has $L_{2}$ error (over Gaussian $x$) at most $\epsilon/(4n)$ for each $m$, then we have that
$\|H(x)-\widetilde{H}(x)\|_{2}$ $\leq\sum_{m=0}^{n}\|{\langle T_{m},H_{m}(x)\rangle}-\widetilde{\langle T_{m},H_{m}(x)\rangle}\|_{2}+\left\|\sum_{m>n}\langle T_{m},H_{m}(x)\rangle\right\|_{2}$
$\leq\sum_{m=0}^{n}\epsilon/(4n)+\sqrt{\sum_{t>n}\|T_{t}\|_{2}^{2}}$
$\leq\epsilon/2+\sqrt{\sum_{t>n}c_{\sigma,t}^{2}\left(\sum_{i=1}^{k}w_{i}\|v_{i}^{\otimes t}\|_{2}\right)^{2}}$
$\leq\epsilon/2+\sqrt{\sum_{t>n}c_{\sigma,t}^{2}}$
$\leq\epsilon.$
In order to compute $\widehat{H}$, we will need use to Proposition 3.2 to approximate $\langle T_{m},H_{m}(x)\rangle$ for all $m\leq n$ for which $c_{\sigma,m}\neq 0$. We do this by noting that $\langle T_{m},H_{m}(x)\rangle=c_{\sigma,m}\langle M_{n},H_{m}(x)\rangle$. In particular, for $t\leq 2m$ even or equal to $m$ we have that
$M_{t}=\mathbf{E}[F(x)H_{t}(x,y)/c_{\sigma,t}]\;,$
where $x$ and $y$ are independent Gaussians. Given samples from $(x,F(x))$ and simulated Gaussians $y$, one can compute $F(x)H_{t}(x,y)/c_{\sigma,t}$ as a sequential tensor computation of order $t$ and size $O(t)$. Similarly, we have that
$H_{m}(x)=\mathbf{E}_{y}[H_{t}(x,y)].$
The second moment of $H_{t}(x,y)/c_{\sigma,t}$ is $O(1)^{t}$ by Lemma 4.3. To bound the second moment of $F(x)H_{t}(x,y)/c_{\sigma,t}$, we note that it suffices to bound
$\mathbf{E}[(F(x)\langle A,H_{t}(x)\rangle/c_{\sigma,t})^{2}]$
for any $t$-tensor $A$ with $\|A\|_{2}\leq 1$. By Lemma 4.3, we have that
$\mathbf{E}[(\langle A,H_{t}(x)\rangle)^{2}]=O(1)^{t}.$
@@
Since $\langle A,H_{t}(x)\rangle$ is a degree-$t$ polynomial of Gaussians, by Gaussian hypercontractivity, we have that
$\mathbf{E}[(\langle A,H_{t}(x)\rangle)^{4}]=O(1)^{t}\;,$
and thus, by Holder’s inequality, that
$\mathbf{E}[(F(x)\langle A,H_{t}(x)\rangle/c_{\sigma,t})^{2}]\leq O(1)^{t}\|F(x)\|_{4}^{2}/c_{\sigma,t}^{2}\leq O(1)^{t}/\delta^{2}\;.$
Thus, for $m\leq n$ and $c_{\sigma,m}\neq 0$, we can apply Proposition 3.2 to get an approximation $\widetilde{\langle T_{m},H_{m}(x)\rangle}$ to ${\langle T_{m},H_{m}(x)\rangle}$ with $L_{2}$ error $\epsilon/(4n)$ with sample and time complexity $\mathrm{poly}(dk/(\epsilon\delta))2^{O(n)}$. This completes our proof. ∎
@@ 4.4 Learning Mixtures of Linear Regressions
Note that the requirement that $\sigma,B\leq 1$ can be removed at the cost of making the runtime $\mathrm{poly}(k,d)(1/\epsilon)^{1+O((B/\sigma)^{2})}$. This is done by computing an upper bound $A$ on $B+\sigma$ (for example by taking many samples and returning something proportional to the largest value of $|y|$), and renormalizing by replacing $y$ by $y/A$.
Proof of Theorem 4.22.
@@ 5 Conclusions
In this work, we gave a general efficient algorithm to approximate higher-order moment tensors, as long as there are reasonable unbiased estimators for these moments that can be efficiently sampled. This type of tensors arise in a range of learning problems. We leveraged our general result to obtain the first polynomial-time algorithms for learning mixtures of linear regressions, learning sums of ReLUs and cosine activations, density estimation for mixtures of spherical Gaussians with bounded means, and parameter estimation for mixtures of spherical Gaussians under optimal separation. In all cases, our learning algorithms run in $\mathrm{poly}(d,k)$ time.